OpenStudy (anonymous):

<1 and <2 are a linear pair. M <1=x-35 and M <2=x+83. Find the measure of each angle. answers - <1=31, <2=149, or <1=66, <2=124

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

m<1 + m<2 = m<3 m<3 = 180 m<1 + m<2 = 180 Substitute the expressions in x for m<1 and m<2 and solve the equation for x. Once you have x, substitute it into m<1 = x - 35 and m<2 = x + 83 to get the measure of the angles.

OpenStudy (anonymous):

I still dont quit get how to do it /:

OpenStudy (mathstudent55):

A linear pair is a pair of angles which have a common side. The non-common sides form a line. The measures of the angles of a linear pair add up to 180 degrees. Since you are given that <1 and <2 are a linear pair, then that means that m<1 + m<2 = 180 You are also given that m<1 = x - 35 and m<2 = x + 83. If you substitute these two expressions in x in m<1 + m<2 = 180, you get the equation: x - 35 + x + 83 = 180 If you solve that equation for x, then you can substitute the value of x in m<1 = x - 35 and m<2 = x + 83 to get the measures of angles 1 and 2.